OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 2x2 + 8y2 = 16. a. Center: (0, 0); Vertices:(0,-2squareroot2)(0,2squareroot2) ; Foci: (0,-squareroot6)(0,squareroot6) b. Center: (0, 0); Vertices: (-8, 0), (8, 0); Foci: (-2squareroot15,0) (2squareroot15,0) c. Center: (0, 0); Vertices: (0, -8), (0, 8); Foci: (0,-2squareroot15),(0,2squareroot15) d. Center: (0, 0); Vertices:(-2squareroot2,0)(2squareroot,0) ; Foci: (-squareroot6,0)(squareroot6,0)

OpenStudy (anonymous):

anyone?

OpenStudy (sidsiddhartha):

opt d

OpenStudy (sidsiddhartha):

eq is:x^2/8+y^2/2=1 a=2sqrt2,b=sqrt2 vertices=(+-a,0)=(+-2sqrt2,0) foci=(+-ae,0) e=sqrt3/2 so foci=(+-sqrt6,0)

OpenStudy (anonymous):

thx so much!